What is the ratio of fundamental frequency of an open end and closed end organ pipe of same length

Grade:11

2 Answers

Khimraj
5 years ago
for open organ pipe fundamental frequency is f1 = v/2l.
for closed organ pipe fundamental frequency is f2 = v/4l.
f1/f2=2
Radha Shukla
5 years ago
Fundamental frequency of open organ pipe=v/2L and that of cop =v/4L...so their ratio will be..2:1... regards radha